A furniture maker has an order for 37 identical tables. If the furniture maker estimates his learning curve at 85% and can reach steady production of 5 hours per table on the 6th table, how long should it take to make all the tables? Express your answer in hours to the nearest whole hour.
Answer #1

Tn=T1*n^b

p=0.85 then b=lnp/ln2=ln(0.85)/ln2=-0.2344

T6=T1*(6^(-0.2344))

T1=7.6103 hours

T2=7.6103*2^(-0.2344) =6.4870 and T3=7.6103*3^(-0.2344)

T4=7.6103*4^(-0.2344);T5=7.6103*5^(-0.2344)

Sum of T1+T2+T3+T4+T5=30.67

Now till 37 the tabke the rate of producing table is 5 hours

32*5=160 hours +30.67 hours

Hence we need 190.67 hours =191 hours to produce identical 37 tables